OpenStudy (anonymous):

The measure of angle A is 70.8º. Find the measures of the complement and supplement of angle A.

OpenStudy (ajprincess):

Sum of complimentary angles=90 Sum of supplementary angles=180

OpenStudy (anonymous):

Complement of angle A = 90 - 70.8 degrees = 19.2 degrees Supplement of angle A = 180 - 70.8 degrees = 109.2 degrees

OpenStudy (anonymous):

thank you